# AsymDex: Leveraging Asymmetry and Relative Motion in Learning Bimanual Dexterity

### v1

@@ Abstract
We present Asymmetric Dexterity (AsymDex), a novel reinforcement learning (RL) framework that can efficiently learn asymmetric bimanual skills for multi-fingered hands without relying on demonstrations, which can be cumbersome to collect. Two crucial ingredients enable AsymDex to reduce the observation and action space dimensions and improve sample efficiency. First, AsymDex leverages the natural asymmetry found in human bimanual manipulation and assigns specific and interdependent roles to each hand: a facilitating hand that moves and reorients the object, and a dominant hand that performs complex manipulations on said object. Second, AsymDex defines and operates over relative observation and action spaces, facilitating responsive coordination between the two hands. Further, AsymDex can be easily integrated with recent advances in grasp learning to handle both the object acquisition phase and the interaction phase of bimanual dexterity. Unlike existing RL-based methods for bimanual dexterity, which are tailored to a specific task, AsymDex can be used to learn a wide variety of bimanual tasks that exhibit asymmetry. Detailed experiments on four simulated asymmetric bimanual dexterous manipulation tasks reveal that AsymDex consistently outperforms strong baselines that challenge its design choices, in terms of success rate and sample efficiency. The project website is at https://sites.google.com/view/asymdex-2024/.
Keywords: Dexterous Manipulation, Bimanual Manipulation
@@ 1 Introduction
We tackle the challenge of learning bimanual dexterous manipulation skills on multi-fingered hands using reinforcement learning (RL). Bimanual skills are crucial for robots operating in human environments as they allow for more complex and flexible manipulation compared to a single hand [1, 2, 3, 4, 5, 6, 7]. While learning dexterous manipulation skills on a single hand presents numerous challenges [8, 9, 10, 11, 12, 13, 14], learning bimanual dexterity can be significantly more challenging due to the higher-dimensional state and action spaces and the need to coordinate and synchronize the movement of two hands [1].
@@
To circumvent the challenges introduced by the high-dimensional state and action spaces of bimanual dexterous manipulation, we take inspiration from how humans approach this challenge. Humans exhibit a natural asymmetry in how we use each of our hands when we perform most bimanual tasks. Specifically, we tend to use one hand to reposition and reorient an object being manipulated so as to make it easier for the other hand to achieve the desired manipulation objectives. While the asymmetric assumption might seem restrictive at first glance, rich bodies of work in human biomechanics and evolution reveal its significance and necessity [15, 16, 17, 18]. Evolutionary biologists posit that such handedness evolved in humans and great apes to meet the escalating cognitive demands of tool use and complex manipulation [19]. Indeed, a large class of real-world bimanual tasks admit this asymmetry (e.g., attachment, detachment, assembly, pouring).
A key insight that we leverage is that the natural asymmetry between the two hands can reduce the dimensionality of the bimanual dexterous manipulation, and in turn improve effectiveness and sample efficiency. Using this insight, we contribute a novel learning framework, dubbed Asymmetric Dexterity (AsymDex), that can efficiently learn asymmetric bimanual dexterous manipulation tasks based on RL (see Fig. 1 for a block diagram). Note that AysmDex does not require the cumbersome collection of demonstrations and can learn bimanual skills only using reinforcement learning.
@@
Figure 1: Our approach (AsymDex) efficiently learns asymmetric bimanual dexterous manipulation skills based on reinforcement learning by effectively leveraging i) the natural asymmetry in the hands’ roles and ii) relative state and action spaces that encourage synchronization.
AsymDex has two crucial ingredients. First, it introduces asymmetry by defining a dominant hand and a facilitating hand. While the facilitating hand learns to reposition and reorient the object, the dominant hand learns complex manipulation skills (including in-hand manipulation). Note that there’s no relative movement between the grasped object and the facilitating hand. As a result, AsymDex holds the fingers of the facilitating hands in the grasping pose and only controls the 6D motion of its base. On the other hand, AsymDex learns to control both the base and fingers of the dominant hand. Second, AsymDex reasons about and controls the relative motion between the dominant and facilitating hands. We define relative observation and action spaces that incentivize responsive coordination between the two hands without resorting to explicit time-dependence.
@@
We also leverage the observation that bimanual manipulation in practice is composed of two distinct phases: i) the acquisition phase in which objects are grasped from surfaces, and ii) the interaction phase where the two hands coordinate to perform the bimanual task. Unlike many existing methods that ignore the acquisition phase, we show that this decomposition enables AysmDex to be seamlessly integrated with learned grasping policies to enable fluent execution.
@@
In summary, we contribute AsymDex – a novel framework for asymmetric bimanual dexterity that leverages the natural asymmetry in hand roles and relative observation and action spaces. We evaluate AsymDex on four complex bimanual dexterous tasks (adapted from BiDexHand [20]) and compare against strong baselines that challenge its design choices. Our results show that AsymDex consistently outperforms the baselines in terms of success rate and sample efficiency.
@@ 2 Related Work
In this section, we contextualize our contributions within relevant sub-fields.
@@
Learning Dexterous Manipulation: Learning dexterous manipulation skills involves addressing numerous challenges due to high dimensional state and action spaces and highly nonlinear dynamics. Recent works have tackled these challenges using imitation learning (IL) or reinforcement learning (RL) and demonstrate impressive performance [8, 9, 34, 10, 35, 11, 12, 36, 37, 14]. However, IL-based methods, including those that combine RL and IL [8, 10], rely either on complex infrastructure and retargeting methods to collect demonstrations  [8, 35, 14, 38, 39] or pre-trained expert policies [11, 12, 37]. On the other hand, RL-based methods do not share these constraints as they learn skills via reinforcement. However, RL methods tend to require single amounts of exploration even for unimanual dexterous manipulation [9, 34, 10, 36]. As we show in our experiments, naive application of RL-based methods is not effective for bimanual dexterous manipulation due to the increased dimensionality and the need for coordination. We also show that AsymDex is able to efficiently learn complex bimanual dexterous manipulation skills using RL without relying on demonstrations.
@@
Learning Bimanual Dexterous Manipulation: A few recent studies have focused on learning bimanual dexterity. Some of these methods require the collection of expert demonstrations [6] and suffer from the same limitations we discussed earlier for IL-based methods that use parallel jaw grippers. To circumvent the need for collecting demonstrations, recent efforts have led to methods that only leverage RL and yet are capable of learning impressive bimanual manipulation skills, such as playing the piano [7], twisting lids off containers [40], and dynamic handover [5]. While these methods are specifically designed to solve a particular task, AsymDex is capable of efficiently learning different bimanual dexterous manipulation tasks.
3 Learning Asymmetric Bimanual Dexterous Manipulation Skills
@@
In this section, we formulate the problem of asymmetric bimanual dexterous manipulation and introduce the different elements of our approach (AsymDex).
3.1 Preliminaries
@@
Consider the problem of bimanual dexterous manipulation, in which two multi-fingered hands coordinate to manipulate up to two objects. Formally, this problem can be defined as a Partially-Observable Markov Decision Process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{Z},\mathcal{A},\mathcal{R},\mathcal{P})$, where $\mathcal{S}\in\mathbb{R}^{n}$ is the state space, $\mathcal{Z}\in\mathbb{R}^{m}$ is the observation space, $\mathcal{A}\in\mathbb{R}^{u}$ is the action space, ${\mathcal{R}:\mathbb{R}^{m}\times\mathbb{R}^{u}\rightarrow\mathbb{R}}$ is the reward function, and ${\mathcal{P}:\mathbb{R}^{n}\times\mathbb{R}^{u}\rightarrow\mathbb{R}^{n}}$ is the environment dynamics. Note that we do not assume access to any demonstrations. Instead, we tackle of challenge of learning purely based on reinforcement. Given this formulation, the problem boils down to learning a policy $\pi:\mathcal{Z}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $E_{\pi}[\Sigma_{t=0}^{T-1}\gamma^{t}\mathcal{R}(z(t),a(t))]$.
@@
Observation and Action Space: The observation space $\mathcal{Z}$ is composed of hand and object measurements. At step $t$, $z(t)$ contains $\xi^{b}_{1}(t)\in\mathbb{R}^{6}$, $\xi^{b}_{2}(t)\in\mathbb{R}^{6}$, $\xi^{h}_{1}(t)\in\mathbb{R}^{n_{1}}$, $\xi^{h}_{2}(t)\in\mathbb{R}^{n_{2}}$, $o_{1}(t)\in\mathbb{R}^{6}$, and $o_{2}(t)\in\mathbb{R}^{6}$, where $\xi^{b}_{1}$ and $\xi^{b}_{2}$ denote the first and the second hand bases’ $6D$ poses, $\xi^{h}_{1}$ and $\xi^{h}_{2}$ denote the first and the second hands’ joint states (e.g., the palm and fingers), and $o_{1}$ and $o_{2}$ represents the $6D$ poses of either two objects (e.g., stacking two cups), or two parts of the same object (e.g., bottle and bottle cap). The action at Step $t$ is given by $a(t)=[\hat{\xi}^{b}_{1}(t),\hat{\xi}^{b}_{2}(t),\hat{\xi}^{h}_{1}(t),\hat{\xi}^{h}_{2}(t)]$, composed of the target base poses ($\hat{\xi}^{b}_{1}$ and $\hat{\xi}^{b}_{2}$) and target joint positions ($\hat{\xi}^{h}_{1}$ and $\hat{\xi}^{h}_{2}$) for both hands. These actions are fed to a PD tracking controller to actuate both hands.
@@
Note that our primary contributions pertain to the interaction phase of bimanual dexterous manipulation, in which two dexterous multi-fingered hands coordinate to complete the task after having grasped the necessary object(s). Most existing works focus solely on the interaction phase [40, 5, 41]. In Section. 3.6, we discuss how our approach can be extended to also tackle the acquisition phase, in which the hands learn to grasp objects before coordinating.
3.2 A Monolithic Approach
We begin by discussing the most straightforward approach one could take: a monolithic policy that utilizes all accessible environment states to plan actions for both hands: $\pi_{\mathrm{naive}}(\hat{\xi}^{b}_{1}(t),\hat{\xi}^{b}_{2}(t),\hat{\xi}^{h}_{1}(t),\hat{\xi}^{h}_{2}(t)|\xi^{b}_{1}(t),\xi^{b}_{2}(t),\xi^{h}_{1}(t),\xi^{h}_{2}(t),o_{1}(t),o_{2}(t))$. Training such a policy can be highly inefficient due to the high dimensionality of the observation and action spaces. Importantly, this monolithic implementation does not exploit the natural asymmetry found in most bimanual tasks. Below, we explain how AsymDex incorporates this insight.
@@ 3.3 Incorporating Asymmetry
When humans execute bimanual manipulation tasks, we tend to use our dominant hand to perform precise manipulations, and use a non-dominant hand to facilitate such manipulation [16, 17, 18]. For instance, when opening a bottle, we typically use our facilitating hand to move and reorient the bottle such that the bottle cap is closer to and oriented toward the dominant hand, which will then grasp the cap and uncap the bottle. During this cooperation process, our non-dominant hand moves and rotates the object with a firm grasp to facilitate the object manipulation by the dominant hand. Motivated by this, we assign different roles to each robot hand (i.e., a facilitating hand and a dominant hand) during bimanual manipulation.
Further, we make the observation that there tends to be no relative motion between the facilitating hand and the grasped object since the facilitating hand need only hold, move, and reorient the object (i.e., no in-hand reorientation). On the contrary, the hand dominant can interact freely with the object either directly or via another object. This observation suggests that the asymmetric manipulation and coordination strategy is neither dependent on nor influences the hand joints of the facilitating hand. As such, we can considerably reduce the observation and action spaces by accounting for asymmetry. Precisely, this observation allows us to define an asymmetric bimanual policy: $\pi_{\mathrm{asym}}(\hat{\xi}^{b}_{d}(t),\hat{\xi}^{b}_{f}(t),\hat{\xi}^{h}_{d}(t)|\xi^{b}_{d}(t),\xi^{b}_{f}(t),\xi^{h}_{d}(t),o_{1}(t),o_{2}(t))$ which use the facilitating hand to only to reposition and reorient the object. This considerable reduction in the dimensions of the observation and action spaces is likely to result in improved sample efficiency.

### v2

# AsymDex: Asymmetry and Relative Coordinates for RL-based Bimanual Dexterity

@@ Abstract
We present Asymmetric Dexterity (AsymDex), a novel and simple reinforcement learning (RL) framework that can efficiently learn a large class of bimanual skills in multi-fingered hands without relying on demonstrations. Two crucial insights enable AsymDex to reduce the observation and action space dimensions and improve sample efficiency. First, true ambidexterity is rare in humans and most of us exhibit strong “handedness”. Inspired by this observation, we assign complementary roles to each hand: the facilitating hand repositions and reorients one object, while the dominant hand performs complex manipulations to achieve the desired result (e.g., opening a bottle cap, or pouring liquids). Second, controlling the relative motion between the hands is crucial for coordination and synchronization of the two hands. As such, we design relative observation and action spaces and leverage a relative-pose tracking controller. Further, we propose a two-phase decomposition in which AsymDex can be readily integrated with recent advances in grasp learning to facilitate both the acquisition and manipulation of objects using two hands. Unlike existing RL-based methods for bimanual dexterity with multi-fingered hands, which are either sample inefficient or tailored to a specific task, AsymDex can efficiently learn a wide variety of bimanual skills that exhibit asymmetry. Detailed experiments on seven asymmetric bimanual dexterous manipulation tasks (four simulated and three real-world) reveal that AsymDex consistently outperforms strong baselines that challenge our design choices. The project website is at https://sites.google.com/view/asymdex-2025/.
Keywords: Dexterous Bimanual Manipulation, Multi-Fingered Hands
@@ 1 Introduction
Bimanual dexterity is crucial for robots operating in human environments as they allow for complex yet flexible manipulation compared to a single hand [1, 2, 3, 4, 5, 6]. We are interested in learning a wide range of bimanual dexterous skills on multi-fingered hands purely from reinforcement.
While learning on a single multi-fingered hand is known to be challenging [7, 8, 9, 10, 11, 12, 13], learning bimanual dexterity is made more challenging due to the higher-dimensionality and the need for coordination and synchronization of two hands [1, 14]. These challenges are only exacerbated when these skills have to be learned from reinforcement, explaining why existing efforts often either resort to expert demonstrations [15, 16, 17, 18] or limit themselves to specific tasks [19, 20].
@@
We rely on two insights to tackle the challenges of bimanual dexterity. First, we are inspired by how humans and other great apes approach these challenges: there is a natural asymmetry in how we use each of our hands when we perform most bimanual tasks [21]. Specifically, we tend to use one hand to reposition and reorient an object so as to make it easier for the other hand to perform complex manipulation. While leveraging such asymmetry might appear to restrict the class of bimanual skills we can learn, rich bodies of work in human biomechanics and evolution reveal its prevalence and necessity [21, 22, 23, 24]. Evolutionary biologists posit handedness evolved to meet the escalating cognitive demands of tool use, language, and complex manipulation [25]. Indeed, a large number of real-world tasks admit such asymmetry (e.g., attachment, detachment, assembly, and pouring).
Looking closely at the asymmetry in bimanual dexterity reveals our second insight. Our non-dominant hand tends to hold an object firmly as we reorient and reposition it relative to the dominant hand or the object being held by the dominant hand (e.g., tilting a pen before uncapping). This suggests that there is often little to no in-hand movement of the object grasped by the non-dominant hand, and robust synchronization can be achieved by ensuring relative movement of the two hands.
@@
Figure 1: Our approach (AsymDex) efficiently learns asymmetric bimanual dexterous manipulation skills based on reinforcement learning by effectively leveraging i) the natural asymmetry in the hands’ roles and ii) relative state and action spaces that prioritize synchronization.
We contribute a novel RL-based learning framework for bimanual dexterity, dubbed Asymmetric Dexterity (AsymDex) by operationalizing the above two insights (see Fig. 3). To incorporate asymmetry, we define a dominant hand and a facilitating hand. While the dominant hand learns complex skills across all its degrees of freedom, the facilitating hand learns to reposition and reorient the object by controlling the 6D pose of its base (i.e., no finger movement). This allows us to both reduce the dimensionality of the observation and action spaces and tightly integrate the roles of the two hands. To ensure coordination and synchronization, AsymDex operates over relative observation and action spaces that incentivize flexible coordination of the two hands without resorting to explicit time-dependence or task-specific coordinate frame designs.
@@
We also leverage the observation that bimanual manipulation in practice is composed of two distinct phases: i) the acquisition phase in which objects are grasped, and ii) the interaction phase in which the two hands coordinate to perform the bimanual task. Unlike many existing methods that entirely ignore the acquisition phase[2, 14, 19, 26], we show that this decomposition enables AysmDex to be seamlessly integrated with learned grasping policies to enable fluent execution.
@@
In summary, we contribute AsymDex – a novel Rl-based framework for learning a wide variety of bimanual dexterous skills by taking inspiration from two key aspects of human bimanual dexterity: i) asymmetric hand roles, and ii) relative hand movement. We conduct comprehensive experiments on seven complex tasks (four simulated and three real) and compare against strong baselines that challenge the need for AsymDex’s structural inductive biases. Our results show that AsymDex consistently outperforms these baselines in terms of both task performance and sample efficiency.
@@ 2 Related Work
Learning Dexterous Manipulation: Learning dexterous manipulation skills involves addressing numerous challenges due to high dimensional state and action spaces and highly nonlinear dynamics. Recent works have tackled these challenges using imitation learning (IL) or reinforcement learning (RL) and demonstrate impressive performance [7, 8, 9, 10, 11, 13, 39, 40, 41, 42]. However, IL-based methods rely either on complex infrastructure and retargeting methods to collect demonstrations  [7, 13, 40, 43, 44] or pre-trained expert policies [10, 11, 42]. On the other hand, RL-based methods do not share these constraints as they learn skills via reinforcement, but tend to require significant exploration even for dexterous manipulation with a single hand [8, 9, 39, 41]. As we show in our experiments, naive application of RL-based methods is not effective for bimanual dexterous manipulation due to the increased dimensionality and the need for coordination.
@@
Learning Bimanual Dexterous Manipulation: A few recent studies have focused on learning bimanual dexterity. Some of these methods require the collection of expert demonstrations [5] and suffer from the same limitations we discussed earlier for IL-based methods that use parallel jaw grippers. To circumvent the need for collecting demonstrations, recent efforts have led to methods that only leverage RL and yet are capable of learning impressive bimanual manipulation skills, such as playing the piano [6], twisting lids off containers [19], and dynamic handover [20]. While these methods are specifically designed to solve a particular task, AsymDex is capable of efficiently learning different bimanual dexterous manipulation tasks. Some recent studies investigate generalized learning method by utilizing expert demonstration for efficient RL training [15, 16, 17, 18], while AsymDex can efficiently learn bimanual manipulation skills without relying on demonstration.
3 Problem Formulation
@@
We first formulate the general problem of bimanual dexterous manipulation, and then introduce asymmetric bimanual dexterity.
@@
Consider the general problem of bimanual dexterous manipulation, in which two multi-fingered hands coordinate to manipulate up to two objects. Formally, this problem can be defined as a Partially-Observable Markov Decision Process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{Z},\mathcal{A},\mathcal{R},\mathcal{P})$, where $\mathcal{S}\in\mathbb{R}^{n}$ is the state space, $\mathcal{Z}\in\mathbb{R}^{m}$ is the observation space, $\mathcal{A}\in\mathbb{R}^{u}$ is the action space, ${\mathcal{R}:\mathbb{R}^{m}\times\mathbb{R}^{u}\rightarrow\mathbb{R}}$ is the reward function, and ${\mathcal{P}:\mathbb{R}^{n}\times\mathbb{R}^{u}\rightarrow\mathbb{R}^{n}}$ is the environment dynamics. Note that we do not assume access to any demonstrations. Instead, we tackle of challenge of learning purely based on reinforcement. Given this formulation, the problem boils down to learning a policy $\pi:\mathcal{Z}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $E_{\pi}[\Sigma_{t=0}^{T-1}\gamma^{t}\mathcal{R}(z(t),a(t))]$.
@@
Observation and Action Spaces: The observation space $\mathcal{Z}$ is composed of hand and object states. At time step $t$, $z(t)=[\xi_{1}(t),\xi_{2}(t),\xi^{obj}(t)]$, where $\xi_{1}(t)$ contains the first hand’s current full (fingers + wrist) configuration $\xi^{h}_{1}(t)\in\mathbb{R}^{n_{1}}$ and the 6 DOF pose of its base $\xi^{b}_{1}(t)\in\mathbb{SE}(3)$, $\xi_{2}(t)$ contains the corresponding elements for the second hand, and $\xi^{obj}(t)$ contains the 6 DOF poses of either two objects (e.g., stacking two cups) or parts of one object (e.g., bottle and lid). The joint action at time step $t$ is given by $a(t)=[\hat{\xi}_{1}(t),\hat{\xi}_{2}(t)]$, where $\hat{\xi}_{1}(t)$ denotes the target joint configuration of the first hand $\hat{\xi}^{h}_{1}(t)$ and the target 6 DOF pose of its base $\hat{\xi}^{b}_{1}(t)$, and $\hat{\xi}_{2}(t)$ denotes the corresponding targets for the second hand. With the above definitions, we can now define the problem of learning bimanual dexterity as one of learning a monolithic symmetric policy: $\pi_{\mathrm{sym}}(a(t)|z(t))$.
Asymmetric Dexterity Problem: Asymmetric dexterity can be viewed as a broad subclass to the above general class of problems. Inspired by strategies employed by humans and other great apes, we are interested in tasks in which one multi-fingered hand performs complex and precise manipulations while the other plays a facilitating role by supporting and actively reorienting objects of interest (e.g., stacking, attachment, detachment, etc.). Note that our formulation does not restrict the movement of the second hand; it merely restricts the relative movement between the second hand and the object being held. As explained in Sec. 1, a large number of bimanual tasks exhibit asymmetry, hinting at handedness in most humans. We are thus interested in learning an asymmetric policy $\pi_{\mathrm{AsymDex}}(\cdot)$ with carefully-defined observation $\mathcal{Z}_{\mathrm{AsymDex}}$ and action $\mathcal{A}_{\mathrm{AsymDex}}$ spaces in an effort to improve both effectiveness and sample efficiency.
@@
Note that our primary contributions and the asymmetric assumption pertain to the interaction phase of bimanual dexterity, in which the two hands actively coordinate to complete the task after having grasped the necessary object(s). Most existing works focus solely on the interaction phase [19, 20, 45]. In Sec. 4.4, we discuss how our approach can be readily extended to also tackle the acquisition phase (learning to grasp the necessary objects before coordinating).
4 AsymDex: Learning Asymmetric Dexterity
While one could learn the symmetric policy $\pi_{sym}(\cdot)$ as defined in Sec. 3, training such a policy can be inefficient or ineffective due to the high dimensionality of the observation and action spaces (see Sec. 5 for empirical evidence). Importantly, a symmetric approach ignores the natural asymmetry found in most bimanual tasks. Below, we explain how AsymDex overcomes these challenges.
@@ 4.1 Incorporating Asymmetry
Motivated by the natural asymmetry in human bimanual manipulation [22, 23, 24], we assign different roles to each robot hand during their interaction: a facilitating hand that is responsible for holding and repositioning and reorienting the object, and a dominant hand which is responsible for fine-grained dexterous manipulation of the object(s). We note that there tends to be no relative motion between the facilitating hand and the grasped object in asymmetric dexterity since the facilitating hand need only hold, move, and reorient the object (i.e., no in-hand reorientation). In contrast, the dominant hand can interact freely with the object(s). This suggests that the asymmetric dexterity is neither dependent on nor influences the finger joints of the facilitating hand during the interaction phase. As such, we can considerably reduce the observation and action spaces by defining an asymmetry-only bimanual policy: $\pi_{\mathrm{asym}}(a_{\mathrm{asym}}(t)|z_{\mathrm{asym}}(t))$ with actions $a_{\mathrm{asym}}(t)=[\hat{\xi}_{d}(t),\hat{\xi}^{b}_{f}(t)]$ and observations $z_{\mathrm{asym}}(t)=[\xi_{d}(t),\xi^{b}_{f}(t),\xi^{obj}(t)]$. This change both reduces the dimensionality and ensures that the facilitating hand doesn’t learn unproductive or unnecessary behaviors.